Ligo Caltech by liaoqinmei

VIEWS: 17 PAGES: 36

									                       Commissioning
                    of the LIGO Detector
                             Rick Savage
                       LIGO Hanford Observatory

                          University of Tokyo
                             July 12, 2002



LIGO-G020286-00-W
                     LIGO Observatories

    HANFORD
    Washington


                                          MIT
                                          Boston




          CALTECH
          Pasadena



                            LIVINGSTON
                            Louisiana



LIGO-G020286-00-W
                    Hanford Observatory (LLO)




          4 km




       2 km


LIGO-G020286-00-W
                    Livingston Observatory (LLO)




          4 km




LIGO-G020286-00-W
                        LIGO Hanford Observatory
                          Commissioning Players
   Scientists Resident at            Grad Students                     Support Staff
    LHO                                 »   Bill Butler – Rochester        »   Dave Barker – CDS
      »   Dick Gustafson – UM           »   Masahiro Ito – UO              »   Greg Mendell – LDAS
      »   Michael Landry – postdoc      »   Rana Rakhola – UO              »   Richard McCarthy –
      »   Fred Raab – Head LHO                                                 Elect.
      »   Rick Savage                 Operators
                                                                           »   Josh Myers – Elect.
      »   Paul Schwinberg               »   Betsy Bland
                                                                           »   Doug Cook – Optics
      »   Daniel Sigg                   »   Corey Gray
                                                                           »   Kyle Ryan – Vac.
   Visiting Scientists                 »   Mark Guenther
                                                                           »   John Worden – Vac.
      »   Stan Whitcomb – CIT           »   Nathan Hindman
                                                                           »   Bartie Rivera –Vac.
      »   Robert Schofield – UO         »   Mark Lubinski
                                                                               prep.
      »   Nergis Mavalvala – MIT        »   Gerardo Moreno
                                                                           »   Christine Patton – Sys.
      »   Peter Fritschel – MIT         »   Hugh Radkins                       Admin.
      »   Rana Adhikari – MIT           »   Cheryl Vorvick                 »   Otto Matherny –
          (grad student)
                                                                               Business
      »   Bill Kells – CIT
                                                                           »   Jill Berry – Admin.
      »   Luca Matone - CIT


    LIGO-G020286-00-W
                             LHO Daily Schedule

   Operators on site seven            Electronic log book (elog) for
    days per week.                      all detector-related activities.
     » Two shifts per day                » Available on the web
             – 7:00 AM to 5:00 PM        http://blue.ligo-wa.caltech.edu/ilog/
             – 4:00 PM to 2:00 AM
                                       Some increased seismic noise
     » Two operators per shift
                                        due to construction of new
   Electronics troubleshooting,        building
    cabling, system maint., etc.         » Typically 6:30 AM to 3:30 PM
    typically in the mornings.
                                       High winds (> 25 mph) usually
   Begin operation of                  make full lock difficult, if not
    interferometers at about 2:00       impossible.
    PM daily.


LIGO-G020286-00-W
               Remote-controlled Interferometer




LIGO-G020286-00-W
                    Hanford Control Room




                             No electronics racks
                             in this room

LIGO-G020286-00-W
                      Epics Control System

-All control via
computer
-Digital length and
alignment control
-Dynamic control
of input matrix in
acquisition mode
(ala M. Evans)



  LIGO-G020286-00-W
                        4k Digital Suspension Control
                                    Screen
   Lots of digital filter
    stages
     » Whitening and
       dewhitening
     » Resonant gain stages
     » Notches
     » Combs
     » Etc.
   Improved optical
    lever performance




    LIGO-G020286-00-W
                    LIGO Control and Data System
                         (D. Barker at LHO)

   12,770 “slow” EPICS               3 MB/sec per ifo. written to
    channels per interferometer        frames
    (LHO 4k ifo. with digital         15 TB of full-frame storage
    suspension controllers)            capacity at LHO
   425 “fast” channels per ifo.      --> Previous 390 hrs. or
    (256 Hz, 2 kHz, 16 kHz)            16.25 days of data available
   60 16-kHz channels per ifo.        on disk
   16 CDS computers (IOCs)           Past 25 days of second-
    per ifo.                           trend data available on disk
                                      All past minute-trend data
                                       available on disk


LIGO-G020286-00-W
                      Automation of Interferometer
                        Configuration Changes
  Common mode transition script
  (D. Sigg)                                                  Automated initial alignment
                                                             (M. Evans)
  # Turn MICH and PRC boost on
  excawrite -b -r "H2:LSC-MICH_SW1R" -l "SET" "H2:LSC-MICH_SW1" 0000000000010000
  ezcawrite -b -r "H2:LSC-PRC_SW1R" -l "SET" "H2:LSC-PRC_SW1" 0000000000010000 sleep 2
  #
  # Turn servo gains to final values
  ezcawrite "H2:LSC-DARM_ERR_K" -1.2
  ezcawrite "H2:LSC-CARM_ERR_K" -1.0
  ezcawrite "H2:LSC-MICH_ERR_K" -0.085
  ezcawrite "H2:LSC-PRC_ERR_K" -0.15 sleep 1 # # Set CM initial state
  ezcawrite -b -r "H2:LSC-CM_SELECT_RD" -l "XOR" "H2:LSC-CM_SELECT" 01000001000000
  ezcawrite H2:LSC-CM_OVERRIDE 0
  ezcawrite H2:LSC-ComMode_CommGainIn 1
  ezcawrite H2:LSC-ComMode_AOGainIn 11
  ezcawrite H2:LSC-ComMode_ALGainIn 3
  ezcawrite H2:IOO-MC_ERR_EXC_ENABLE 0
  ezcawrite H2:IOO-MC3_LSC_IN 1
  ezcawrite H2:LSC-ComMode_OpenLoop 1
  #




LIGO-G020286-00-W
                           “One-click” Generation of
                         Calibrated Spectra (M. Landry)
– uses most recent
  actuation
  calibrations              DTT software package
– makes swept sine          by D. Sigg
  measurement of
  transfer function
  from TM drive
  toAS_Q
– pole-zero fit using
  Root
– measures AS_Q
  spectrum and
  applies calibration



     LIGO-G020286-00-W
                    LHO 2k Interferometer Status

   5 Watts laser power                      Tidal prediction applied to
    delivered to modecleaner.                 reference cavity
     » Attenuation at AS port gives 20-       temperature.
       40 mW effective power
   Recycling factor 15-25                   Fine actuators (PZTs on
     » Contaminated optic?                    ETM stacks) used for
   Analog suspension                         residual tidal compensation.
    controllers.                             First part of intensity
   Only WFS 1 active                         stabilization loop is active.
     » Improved optical lever damping        Automated calibration.
       for ETMs
   Common-mode servo                        Multi-hour locked stretches,
    configuration changed (now                often 6-10 hours.
    similar to TAMA)

LIGO-G020286-00-W
                    Sensitivity Improvement
                        Hanford 2k Ifo.




LIGO-G020286-00-W
                    LHO 4k Interferometer Status

   1 Watt laser power delivered           Tidal prediction to reference
    to modecleaner                          cavity temperature
     » 20 mW effective at AS port
                                           Fine actuators on ETM
   Recycling factor 40-50
                                            stacks for balance of tidal.
   Digital suspension
    controllers                            Inner loop of intensity
     » Lots of digital filter stages        stabilization active.
     » Frequency-dependent                 Lots of code changes.
       diagonalization of drive of TM
       drives                              Shakedown of digital
   Only WF1 active                         suspension code now
   Original common-mode                    (almost) complete.
    servo configuration                    Now locking for up to 8
                                            hours.

LIGO-G020286-00-W
                    Sensitivity Improvement
                       Hanford 4K Ifo.




LIGO-G020286-00-W
                    LLO 4k Interferometer Status

   1.9 Watt laser power to MC.      Feed-forward to ETMs to
   Recycling gain ~ 50.              reduce large microseism.
   Analog suspension                Inner loop of intensity servo
    controllers.                      active.
   Original common-mode             Unable to lock during logging
    servo configuration.              – most weekdays.
   Closed-loop control of ref.      Adversely affected by traffic
    cavity temp. for CM tides.        on highway.
   Fine actuators for residual      Lock broken by trains twice
    tidal correction.                 nightly.
   PEPI to reduce 0.3 –5 Hz         Locking for multi-hour
    seismic noise at ETMs             stretches when noise allows.

LIGO-G020286-00-W
                    Sensitivity Improvement
                       Livingston 4k Ifo.




LIGO-G020286-00-W
                    Interferometer noise modeling
                             (R. Adhikari)
                                          EO shutter at AS
                                           port attenuates light
                                           by factor of 50
                                          New optical lever
                                           whitening filters
                                           should reduce ADC
                                           noise by factor of
                                           400
                                          Below 15 Hz, freq.
                                           noise resulting from
                                           MC radiation
                                           pressure fluctuations
                                           dominates



LIGO-G020286-00-W
                        2k Sensitivity Improvement:
                        Reconfiguration of CM servo
     Original configuration of frequency stabilization topology




                                                From Nergis M.


    LIGO-G020286-00-W
                 Reconfiguration of Common-mode
                         Servo (D. Sigg)
     Eliminate CARM drive to ETMs
     MC servo feedback to laser frequency actuator only




                                             From Nergis M.
    LIGO-G020286-00-W
              Improved Displacement Spectrum




LIGO-G020286-00-W
                    Relative Intensity Noise
                          Stabilization
                                    4k laser current shunt xfer. fctn.




   Free-running slow power
    variations ~1-2%
   Insufficient drive range,   D.C level = 1.7 V  0.37%/V
    causing power glitches            Design range > +,- 2.5%
    (with Y. Aso, 5/02)
LIGO-G020286-00-W
                     New Intensity Stabilization Servo

   New two-loop servo
                              LHO 4k Outside-the-loop RIN Spectra
    and new low-noise
    photodetectors
   Inner loop with PD just
    after pre-modecleaner
   Outer loop with PD just
                                                      Free-running
    after modecleaner
    (outside vacuum)
   RIN goal: 1e-8/rtHz
   Only inner loop
    installed, June, 2002
   +,- 10 V drive range                                Stabilized
DP/P = 5e-8 Hz-1/2


     LIGO-G020286-00-W
                    LHO 4k RIN Spectra

Gain ~ 30 dB at 1 Hz




                          RIN not yet measured downstream
                          of MC with this servo.

LIGO-G020286-00-W
                    Grounding and EMI
                              Frequency stabilization
60 Hz and harmonics            servo error point
                                » Green – laser power supply in
                                  rack
                                » Red – laser power supply
                                  moved to top of PSL enclosure




LIGO-G020286-00-W
                    Parasitic Interferometer
                            LHO 2k
                                  Measurement made with RM
                                   aligned and other optics
                                   misaligned
                                  1 Hz period indicates small
                                   optic motion
                                  6 fringes >> 3 micronspeak
                                  Tried to identify which optic
                                   moving by changing
                                   damping gains – no change
                                   observed
                                  Investigaion abandoned
                                   because edge is below
                                   seismic wall


LIGO-G020286-00-W
              Source of Parasitic Interferometer
                         Identified?
                    Non-stationary noise      Post MMT2 slippage
                    from parasitic ifo.        investigation
                                               » “uncovered an unsettling
                                                 scenario. From the plots, SM1
                                                 has been without it's
                                                 photodiode sensor readbacks
                                                 (and damping) since
                    Measured downstream          ~12/18/01.” J. Myers July1, 2002 elog
                    of modecleaner                entry



                                                          electronics module
                                                          failed 12/18/01




LIGO-G020286-00-W
                                 Delay of S1 Run
                              Friday, June 28, 2002

                    2k mode-matching telescope small optic
                    pitch monitor channel
                                  30 minutes
                                                      optic drops

                                               slip? starts

                      LHO x-end vertical seismometer channel
                                               peak to peak vertical velocity in
                                               the spike is 80 microns/sec




LIGO-G020286-00-W
                    Slack MMT2 Suspension Wire


                                     Suspension wire




LIGO-G020286-00-W
                    Repair of MMT2 Suspension
“A preliminary inspection of mirror        Working hypothesis is that
    MMT2 shows two broken
                                            the loss of control of one of
    magnet assemblies. The wire
    break is along the free section         the other optics (SM2)
    of wire, not in the clamp. The          caused the laser beam to hit
    optic is on its way to the optics       the suspension wire of
    lab for a post-mortem and               MMT2, precipitating the
    repairs. Chamber door is going          breakage.
    back on and we hope to be              Tests of breaking wires
    pumping soon.”
                                            under load with hot irons
                     Fred Raab7/2/02        appear to support this
   Earthquake stops did not                hypothesis.
    prevent magnet breakage –              New baffles to protect
    re-design underway.                     suspension wires designed
                                            and implemented.
LIGO-G020286-00-W
               New Baffles to Protect Suspension
                   Wires from Laser Beams

           MMT2                           MC2


New baffle                               Baffle
                                         extension


                                         Black glass
 Alignment                               baffle
 fixture


  LIGO-G020286-00-W
                                       Pumpdown
   MMT2 repaired and re-
                                                       Waiting for water vapor
    installed
                                                        pressure to decrease
   Pumpdown began at 12:35                             so that gate valves to
    PST on July 9.                                      beam tubes can be
                                                        opened.
                        Fit to early                    Preliminary alignment,
                        pumpdown                        centering on Faraday
                        trend                           isolator, etc., looks
                                                        good.


                                        1e-7 torr      It looks like the S1 run
                                       on July 18       will begin the day after
                                                        the LSC meeting on
                                                        August 23, 2002.
    LIGO-G020286-00-W
                               After the S1 Run
     Still lots to be done!          Many pending hardware and
                                       software installation tasks




                                                         P. Fritschel 6/02




    LIGO-G020286-00-W
                                LIGO Run Plan

   Science 1 run – Upper Limits            “During 2003 and 2004,
     » June 29 to July 15 (Delayed)          we will plan to run in this
     » 2.5 weeks comparable to E7            search mode for at least
     » Target sensitivity: 200x design       50% of the calendar time,
   Science 2 run – Upper Limits             followed by the planned
     » Nov. 22 to Jan. 6, 2003               one year integrated LIGO
     » 8 weeks                               science run at design
     » Target sensitivity: 20x design        sensitivity. This science
   Science 3 run – Search Run               run will be completed prior
     » July 1, 2003 to Jan. 1, 2004          to proposed major
     » 26 weeks                              interferometer
     » Target sensitivity: 5x design         replacements.”
                                                 LIGO Lab Planning Memo.


LIGO-G020286-00-W

								
To top